Nico is saving money for his college education. He invests some money at 9%, and $1400 less than that amount at 4%. The investments produced a total of $217 interest on 1 year. How much did he invest at each rate?
Answer by dkppathak(439)

Nico is saving money for his college education. He invests some money at 9%, and $1400 less than that amount at 4%. The investments produced a total of $217 interest on 1 year. How much did he invest at each rate?
let he invest$ X at 9% and $X-1400 at 4%
interest on first= 9x/100
interest on second =4(x-1400)/100
9x/100 +4(x-1400)/100 =$217
9x+4x-5600=21700
13x=21700+5600
13x=27300
x=27300/13
=$2100 and other will be 2100-1400=$700